OpenStudy (anonymous):

Use factoring to find the zeros y=x^2-7x-8 and the axis of syymetry .

OpenStudy (anonymous):

PUT Y=O x^2-8x+x-8=0 x(x-8)+1(x-8)=0 (x+1)(x-8)=0 x=-1 or x=8 noe the axis of symmetry x=7/2